OpenStudy (anonymous):

What is the period of x(t)=6cos((2*pi*t )/7)+2cos((4pi*t)/7)

OpenStudy (anonymous):

Why is it 7?

OpenStudy (tkhunny):

\(\dfrac{2\pi}{\dfrac{2\pi}{7}} = 7\) \(\dfrac{2\pi}{\dfrac{4\pi}{7}} = 7/2\) Which is longer?